OpenStudy (anonymous):

An airplane flew 3.5 hours with a 40 mph head wind. The return trip with a tail wind of the same speed took 2.5 hours. Find the speed of the plane in still air.

OpenStudy (mani_jha):

Let the speed of the plane in still air be v. Speed in the first journey = v+40 Speed in the return journey= v-40 Let the distance be x. Speed = Distance/Time You think you can do this now?

OpenStudy (anonymous):

lets see...

OpenStudy (anonymous):

\[v+40=x/2.5\] \[v-40=x/3.5\]

OpenStudy (mani_jha):

v+40=x/3.5(This is for the first journey) v-40=x/2.5(This is for the return journey)

OpenStudy (anonymous):

v is the speed in still air?

OpenStudy (mani_jha):

Yes. Now, can you solve these two equations?

OpenStudy (anonymous):

i used elimination and got v=x/3.5 and v=x/2.5

OpenStudy (anonymous):

ok i understand where u r comin from

OpenStudy (mani_jha):

@Luis Rivera, shouldnt it be 3.5x-140=2.5x-100 You wrote a + before 100. @ariesdragon20, follow the method shown by Luis. Did u understand it?

OpenStudy (anonymous):

yeah thanks a lot both of you

OpenStudy (mani_jha):

Oh ya, you are right. Sorry